victoriag:

The equation of line CD is y = 3x − 3. Write an equation of a line perpendicular to line CD in slope-intercept form that contains point (3, 1).

snowflake0531:

Step one: Find the slope of the perpendicular line Step two: Use the slope and the coordinate point, and equation, find y intercept Perpendicular slope is negative reciprocal, so the slope would be \(-\frac{1}{3}\) Then use this and the coordinate and put it into the equation \(1=-\frac{1}{3}(3)+b\\1=-1+b\\b=2\) So the equation would be \(y=-\frac{1}{3}+2\)